# A Dual-Motor Actuator for Ceiling Robots with High Force and High Speed Capabilities

### v1

@@ V EXPERIMENTS
To better comprehend the functionality of the system, Figure 9 shows a complete trial that includes all three modes of operation (patient assistance, fall prevention, and patient transfer).
@@ V-A Force fidelity
Experiments were done to compare the different patient assistance control algorithms. The quality of force is measured with eight healthy adults [^1] All tests were approved at Université de Sherbrooke by the Comité d’éthique de la recherche – Lettres et sciences humaines on Dec. 13 2022 on a predetermined course (see Figure 10). The user starts from a sitting position, stands up, walks for four meters, and sits back down, similarly to the “get-up and go” [24] test, while the system unloads 200 N which was determined as a comfortable unloading of force to the users. The experiment is repeated with the different control algorithms. The force applied to the output is measured using the strap’s load cell during experiments. Mean absolute error is calculated and compared for sit-to-stand and walking phases. The speed is also measured using the encoders on the motors to separate the walking and the sit-to-stand phases. During the sit-to-stand phase, the system moves at higher speeds, above 0.3 m/s. For the walking phase, the system’s speed is lower than 0.3 m/s (see Table III).
@@
TABLE III: Absolute mean error of force (N) for the different algorithms
@@ V-B Fall prevention and recovery
To validate the performances of the brake during fall prevention, tests with different weights (68 kg, 90 kg, 113 kg) were conducted with a desired deceleration of $1~{}m/s^{2}$, $2~{}m/s^{2}$ and the maximum possible braking force to find the minimal fall distance. Figure 11 shows the fall sequence of a test with a 90 kg weight at $2~{}m/s^{2}$. a) Starting at a height of 1 m, the weight gradually falls until, b) reaching 90 cm/s, triggering the system to engage the brake for controlled deceleration. To help downshift faster, EM1 applies its maximum downward velocity (see equation 1). c) Once EM2 stops, EM1 takes over to slow the fall until, d) it stops completely. e) Finally, EM1 lifts the patient to the height at which the fall was detected to facilitate the patient’s ability to stand up and resume their activity.
Figure 11: Complete fall sequence with the output speed, the speed contribution of EM2 and of EM1 (see equation. 1), the output force, the servo angle, and the output position.
By measuring the height at the detection of the fall and at the lowest point, it is possible to find the fall distance $x$. The average acceleration during the fall is calculated with $a=v_{i}^{2}/(2x)$ with $v_{i}$ being the velocity of the mass at the detection of the fall, and the average force is found with $F=m(g+a)$. Table IV shows the results of the tests conducted. Measures for the average acceleration, fall distance and average forces were compared to their theoretical values. The acceleration was the parameter that was adjusted and is related to the patient’s comfort. The fall distance is more intuitive to understand and is linked to the patient’s safety since it indicates the minimal falling distance that the patient needs. The force roughly corresponds to the braking capabilities of the system.
TABLE IV: Fall comparison tests
@@ VI CONCLUSIONS
This paper introduces a prototype ceiling robot capable of patient transfer, assistance, and fall prevention using a dual-motor actuator. The prototype lifts patients up to 318 kg at 0.05 m/s in transfer mode and unloads up to 59 kg at 0.55 m/s in assistance mode. During patient assistance, force tracking error was 12% at high speed and 7.8% at low speed using a friction compensation algorithm and both motors, comparable to a traditional closed-loop controller. For the fall prevention algorithm, it is possible to control the deceleration of the patient by using the disk brake to apply a desired amount of force to dissipate energy.
Future work involves clinical evaluation to validate if the technology meets patient needs. This could determine if the force controller needs further improvement for patient comfort. To better control the braking of a fall and ensure the safety of all patients, the servo motor actuator could be replaced by a faster and stronger brake. Also, the brake could be normally closed to increase the safety of the patients. While suitable for patient lifts, the actuators could also serve a wider range of robots with dual operation points and rapid downshifting capabilities.

### v2

@@ V EXPERIMENTS
To better comprehend the functionality of the system, Figure 9 shows a complete trial that includes all three modes of operation (patient assistance, fall prevention, and patient transfer). Trials for patient assistance and fall prevention are described in more details in the following paragraphs.
@@ V-A Force fidelity
Experiments were done to compare the different patient assistance control algorithms. The quality of force is measured with eight healthy adults [^1] All tests were approved at Université de Sherbrooke by the Comité d’éthique de la recherche – Lettres et sciences humaines on Dec. 13 2022 (age: from 22 to 29 years old, weight: from 45 to 100 kg, 2 females and 6 males) on a predetermined course (see Figure 10). The user starts from a sitting position, stands up, walks for four meters, and sits back down, similarly to the “get-up and go” [30] test, while the system unloads 200 N which was determined as a comfortable unloading of force to the users. The experiment is repeated with the different control algorithms. The force applied to the output is measured using the strap’s load cell during experiments. Mean absolute error is calculated and compared for standing up, sitting down, and walking phases. The speed is also measured using the encoders on the motors to separate the walking and the standing/sitting phases. The system moves at higher speeds, above 0.3 m/s during the standing and sitting phases. For the walking phase, the system’s speed is lower than 0.3 m/s (see Table IV). The friction compensation with EM1 algorithm is then compared to every other algorithm using a Mann-Whitney test.
@@
TABLE IV: Absolute mean error of force (N) for the different algorithms of Figure 7
@@ V-B Fall prevention and recovery
To validate the performances of the brake during fall prevention, tests with different weights (68 kg, 90 kg, 113 kg) were conducted with a desired deceleration of $1~{}m/s^{2}$, $2~{}m/s^{2}$ and the maximum possible braking force to find the minimal fall distance. Figure 11 shows the fall sequence of a test with a 90 kg weight at $2~{}m/s^{2}$. a) Starting at a height of 1 m, the weight gradually falls until, b) the encoder in EM2 reads an output speed of 0.90 m/s, triggering the system to engage the brake for controlled deceleration. To help downshift faster, EM1 applies its maximum downward velocity (see equation 1). c) Once EM2 stops, EM1 takes over to slow the fall until, d) it stops completely. e) Finally, since only EM1 can lift the full weight of the patient, EM1 lifts the patient to the height at which the fall was detected to facilitate the patient’s ability to stand up and resume their activity.
Figure 11: Complete fall sequence with the output speed, the speed contribution of EM2 and of EM1 (see equation. 1), the output force, the servo angle, and the output position. The different events are: a) start of the fall, b) start of the braking sequence, c) EM2 stops, d) EM1 stops, e) end of patient recovery.
By measuring the height at the detection of the fall and at the lowest point, it is possible to find the fall distance $x$. The average acceleration during the fall is calculated with $a=v_{i}^{2}/(2x)$ with $v_{i}$ being the velocity of the mass at the detection of the fall, and the average force is found with $F=m(g+a)$. Table V shows the results of the tests conducted. Measures for the average acceleration, fall distance and average forces were compared to their theoretical values. The acceleration was the parameter that was adjusted and is related to the patient’s comfort. The fall distance is more intuitive to understand and is linked to the patient’s safety since it indicates the minimal falling distance that the patient needs. The force roughly corresponds to the braking capabilities of the system.
TABLE V: Fall comparison tests
@@ VI CONCLUSIONS
This paper introduces a prototype ceiling robot capable of patient transfer, assistance, and fall prevention using a dual-motor actuator. The prototype lifts patients up to 318 kg at 0.05 m/s in transfer mode and unloads up to 59 kg at 0.55 m/s in assistance mode. During patient assistance, force tracking error was 12% at high speed and 7.8% at low speed using a friction compensation algorithm and both motors, comparable to a traditional closed-loop controller. This is over the 5% error target, but it was only tested with eight healthy adults as a preliminary validation, and no participant mentioned discomfort during the tests. Future work involves clinical evaluation with patients of different profiles to validate if the technology meets patient needs. This could determine if the force controller needs further improvement for patient comfort.
For the fall prevention algorithm, the patient’s deceleration can be controlled by using the disk brake to apply a desired amount of force to dissipate energy. To better control the braking of a fall and ensure the safety of all patients, the servo motor actuator could be replaced by a faster and stronger brake. Also, the brake could be normally closed to increase the safety of the patients. While suitable for patient lifts, the actuators could also serve a wider range of robots with dual operation points and rapid downshifting capabilities.